# LLM4WM: Adapting LLM for Wireless Multi-Tasking

### v1

@@ I Introduction
The wireless channel plays a critical role in determining the quality and reliability of communication. To achieve the low latency and high reliability required for these applications, millimeter-wave (mmWave) and Multiple-Input Multiple-Output (MIMO) technologies are among the most promising solutions [1, 2, 3]. Accurate channel estimation is vital for maximizing the benefits of these technologies [4]. However, this task becomes increasingly challenging as the number of antennas grows and environmental dynamics become more complex. Understanding the wireless channel’s characteristics—such as fading, interference, and multipath propagation—is crucial for optimizing performance. Fortunately, artificial intelligence (AI) has significantly improved channel estimation accuracy, fostering optimism for the practical implementation of advanced wireless applications [5, 6, 7]. AI also enhances various communication tasks, including channel prediction, beamforming, and positioning, demonstrating notable effectiveness and robustness.
@@
Recently, large language models (LLMs) have emerged as powerful multi-task learners, demonstrating remarkable reasoning capabilities and generalization across various domains, including natural language processing (NLP) [13], healthcare [14], law [15], and finance [16]. The GPT-4 has shown outstanding performance in NLP tasks, while models like BeiT [17] have achieved state-of-the-art results in computer vision tasks such as object detection and semantic segmentation. Inspired by these successes, there is growing interest in leveraging pre-trained models for cross-domain tasks, including channel-associated tasks in wireless communication. For instance, an LLM-empowered channel prediction method LLM4CP was proposed in [18], achieving greatly improved few-shot generalization ability. A foundational channel model WiFo [19] was trained on a diverse channel dataset to perform tasks like time-domain and frequency-domain prediction with zero-shot learning, but they primarily focus on channel reconstruction. Our objective is to enhance multiple channel-associated tasks in wireless communication using LLMs, which poses challenges such as effectively transferring cross-domain knowledge and managing task diversity. We address these challenges through the MoE-LoRA fine-tuning method [20], designed for multi-task scenarios, while noting that recent work has not sufficiently considered the specific relationships between channel-associated tasks, potentially hindering the model’s ability to learn general representations.
Specifically, we propose a multi-task fine-tuning framework based on pre-trained large language models, which is successfully applied to the channel-associated multi-tasks. Unlike previous multi-task learning approaches that rely on shared bottom layers, we freeze most of the large model’s parameters and introduce MoE-LoRA for multi-task fine-tuning. On the one hand, tasks can share experts’ weight, which helps the network learn common knowledge across tasks, while on the other hand, the independence of experts and the gating mechanism ensure the differentiation of task-specific features. Additionally, we incorporate multi-task adapter modules at both the input and output layers of the LLM to ensure alignment between the feature space of communication tasks and the semantic feature space of the model. To further enhance our framework, we also design a pre-processing method and corresponding output header for each task, which optimizes the overall performance and adaptability of the model. The key contributions of our work are summarized as follows:
@@
  • The model demonstrates strong performance across a variety of tasks, including channel estimation, channel prediction, localization enhancement, and beam management. Furthermore, it exhibits impressive generalization capabilities, highlighting robustness and versatility for diverse applications in the wireless domain.
@@ II SYSTEM DESCRIPTION
A dual-frequency communication system operating at sub-6G and mmWave frequencies is considered, which consists of one base station (BS) and one user equipment (UE). Both the BS and UE utilize multiple-input multiple-output (MISO) and orthogonal frequency division multiplexing (OFDM) technologies, with each equipped with transceivers for both frequency bands. The sub-6G and mmWave antennas are co-located and aligned, with similar apertures, enabling them to share spatial features. For clarity, we use $\tilde{\left(\cdot\right)}$ notation, such as $\tilde{x}$ to denote parameters related to the sub-6G system. In mmWave band, the BS adopts an analog beamforming architecture and is equipped with $N_{t}$ uniform linear array (ULA) antennas, while in the sub-6G band, it utilizes a fully digital beamforming architecture with $\tilde{N}_{t}$ ULA antennas. The UE is equipped with an omnidirectional antenna for both frequency links and can accommodate multiple antennas through parallel processing. While the analysis assumes line-of-sight (LoS) conditions for both low-frequency and mmWave channels, the proposed scheme can be easily adapted for scenarios with multiple dominant paths.
@@ II-A Channel Model
In this context, $N$ and $M_{n}$ are the number of clusters and paths in each cluster, respectively. $\beta_{n,m}$, $\upsilon_{n,m}$, $\tau_{n,m}$, and $\Phi_{n,m}$ represent the complex path gain, doppler frequency shift, delay, and random phase, respectively. $\bm{a}(\theta_{n,m})$ represents the steering vector of the corresponding path, where $\theta_{n,m}$ denote the azimuth angles of departure (AoD). Considering the structural characteristics of ULA, the expression for $\bm{a}(\theta_{n,m})$ is derived as:
@@ II-B Signal Model
where $\bm{n}_{t,k}$ is the additive white gaussian noise (AWGN) and $\bm{w}_{t}\in\mathbb{R}^{{N}_{v}\times 1}$ is the beam vector which is selected from predefined codebooks ${\bm{\mathcal{W}}}$, i.e. $\bm{w}_{t}\in{\bm{\mathcal{W}}}$ The achievable spectral efficiency (SE) [21] of the downlink transmission process is derived as:
@@
Similarly, the sub-6G downlink signal transmission process also adopts MISO-OFDM, and it has the same transmission expression as the mmWave link, which will be omitted here for brevity. However, the sub-6G link employs digital precoding, so the LS method estimates the channel at the pilot positions to get the uplink channel as follows:
@@
where $\tilde{w}_{t,k}^{*}$ represents the beam vectors of time $t$ and the $-k\mbox{-}$th subcarrier at sub-6G link. It is worth noting that the effectiveness of $\tilde{w}_{t,k}^{*}$ depends on the accuracy of $\tilde{{h}}_{LS}(t,f_{k})$. Specifically, inaccurate $\tilde{{h}}_{LS}(t,f_{k})$ will lead to mismatched $\tilde{w}_{t,k}^{*}$ with the actual $\tilde{h}(t,f_{k})$, reducing the received signal-to-noise ratio (SNR) and thereby impairing the SE. Therefore, accurate CSI is crucial for improving the SE of the system.
@@ III-A Channel Reconstruction: Interpolation and Prediction
The objective of channel reconstruction tasks is typically to utilize known channel matrices for the prediction or interpolation of the target channel matrix in the time-frequency domain. The model should learn the inter-domain correlations of the channel matrix across the time, frequency, and antenna domains. So the channel matrix $\bm{H}$ considered herein includes three dimensions: time, space, and antenna, and can be expressed as follows:
@@
For the channel estimation task, a comb-type pilot pattern is utilized, featuring continuous pilots in the time-domain resource blocks (RBs) and discrete distribution in the frequency-domain RBs. The network must learn channel characteristics across various frequency points and fill in the missing channels at absent frequency points. Consequently, $\bm{X}^{I}_{CE}$ and $\bm{X}^{L}_{CE}$ are defined as follows:
@@
Here, $\tilde{T}$ and $\tilde{K}$ denote the total number of timestamps and the number of OFDM symbols. $n_{pilot}$ denotes the number of pilots, which is typically set such that $\tilde{K}/n_{pilot}=4$.
@@ III-B Beam Management: Sub-6G Aided Beamforming
Beamforming requires accurate acquisition of the optimal weight vector $\mathbf{w}_{t}$ based on the codebook $\mathbf{W}\in\mathbb{R}^{N_{v}\times N_{c}}$, where $N_{c}$ is the codebook size. To achieve higher spatial resolution, a super-resolution Discrete Fourier Transform (DFT) codebook is applied, which implies that $N_{c}>N_{v}$. Given that the main paths in sub-6G and mmWave bands are often highly correlated in the LoS conditions, utilizing sub-6 GHz to assist mmWave beamforming is highly meaningful. The model extracts the main path information from the sub-6 GHz channel matrix and serves mmWave beamforming, thus the $I_{BF}$ and $L_{BF}$ for sub-6G aided beamforming task can be defined as:
@@ III-C Radio Environment Mining: Distance Estimation and Path Loss Estimation
Due to the rich environmental information contained in the acquired channel data, such as the multipath components reflecting the density of surrounding buildings, and the delay indicating the distance to the target, the task of radio environment mining aims to leverage the estimated channel to extract environmental information, such as the distance $x_{d}$ between the UE and the BS and the path loss $x_{pl}$ of the main path. This extracted information can be utilized to adjust the configuration of the communication system, ultimately improving communication quality. Thus, $\bm{X}^{I}_{DE}$ and $\bm{X}^{L}_{DE}$ for distance estimation task can be defined as:
@@ IV LLM FOR Wireless Channel-Associated Tasks
All of these tasks mentioned above are fundamentally rooted in the wireless channel, suggesting that multi-task learning can significantly enhance the model’s ability to extract generalizable channel representations. This, in turn, would improve the performance of each individual task. To leverage this potential, we propose a multi-task method for wireless channel applications, empowered by LLM, which we refer to as LLM4WM. This approach integrates the learning of these tasks within a unified network framework based on LLM. Below, we provide a detailed overview of the network components and the training process for LLM4WM.
@@ IV-A Preprocessor Module
where the $\text{Flatten}(\bm{X},i)$ operation denotes flattening the $i$-th dimension of the tensor $\bm{X}$ and all subsequent dimensions into a single dimension. For tasks such as beamforming, distance estimation, and path loss estimation that require channel angle features, the CSI data will undergo domain transformation to convert the spatial domain CSI to angle domain CSI, i.e.,
@@ IV-B Multi-Task Adapter Module
We extend the conventional use of adapter modules, which are small, trainable layers inserted between the layers of a pre-trained language model [27, 28]. This approach allows the model to maintain its existing modeling and generalization capabilities while adapting to a specific domain. However, it is primarily designed for single-task scenarios and lacks the ability to facilitate transfer generalization across multiple tasks simultaneously.
@@
where $L$ and $D_{\text{llm}}$ denote the token length of the input to the LLM and the hidden layer dimension of the LLM, respectively. Since the preprocessed features are two-dimensional data, the $\text{Linear}(\cdot)$ operation includes at least two fully connected operations, which linearly map the first and second dimensions to the specified dimensions. Then the residual feature extraction networks and the activation function would act on $\bm{X}_{n}^{f}$ to obtain feature maps with aligned semantic features as:
@@ IV-C Mixture-of-LoRA Based Fine-tuning
Assume the pre-trained weights are $W_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, where $d_{\text{in}}$ is the input dimension and $d_{\text{out}}$ is the output dimension. The two trainable low-rank matrices are $A\in\mathbb{R}^{r\times d_{\text{in}}}$ and $B\in\mathbb{R}^{d_{\text{out}}\times r}$. Then, we can obtain the fine-tune weights $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ as follows:
@@
where $r$ is the rank of the low-rank approximation. The hyperparameter $\alpha$ allows for adjusting the rank $r$ and is typically set as $\alpha=2\times r$.
Assuming the input to the feed-forward network is $x_{t}$ and the output is $h_{t}$, the forward propagation of the model can be expressed as follows:
$h_{t}$ $=Wx_{t}=W_{0}x_{t}+\frac{\alpha}{r}BAx_{t}.$ (23)
@@
$h_{t}=Wx_{t}=W_{0}x_{t}+\frac{\alpha}{r}\sum_{k=1}^{N_{e}}{\omega}_{k}B_{k}A_{k}x_{t},$ (24)
@@
Notably, the design of the gating network directly affects the performance of the MoE model. To avoid overfitting, we use a single-layer linear network to generate the expert weights for each task, and normalize the weight matrix using the $\text{Softmax}(\cdot)$ function to maintain the stability of the output data.
We apply MoE-LoRA to the linear layers within the feedforward network (FFN) of LLM, while keeping the rest of the parameters frozen. This approach significantly reduces the number of tunable parameters in the model, greatly lowering training costs and improving training efficiency.
@@ IV-D Multi-Task Output Module
To align the task output with the LLM’s semantic feature space, we use a multi-task adapter connected directly to the LLM’s output, which is identical to the input adapter. Assuming the task $n$’s output feature of the large model is $\bm{X}_{n}^{LLM}$, and the multi-task adapter for the task $n$ is denoted as $\text{Adapter}_{n}^{out}$, so this process can be represented as:

### v2

@@ I Introduction
The quality and reliability of communication are largely determined by the wireless channel, which plays a crucial role in this process. To achieve low latency and high reliability, millimeter-wave (mmWave) and Multiple-Input Multiple-Output (MIMO) technologies are among the most promising solutions [1, 2, 3]. Maximizing the benefits of these technologies relies heavily on accurate channel estimation [4]. However, this task becomes increasingly challenging as the number of antennas grows and environmental dynamics become more complex. Understanding the wireless channel’s characteristics—such as fading, interference, and multipath propagation—is crucial for optimizing communication performance, and also supports other related technologies, such as the design of integrated sensing and communications (ISAC) systems [5]. Fortunately, artificial intelligence (AI) has significantly improved channel estimation accuracy, fostering optimism for the practical implementation of advanced wireless applications [6, 7, 8]. AI also enhances various communication tasks, including channel prediction, beamforming, and positioning, demonstrating notable effectiveness and robustness.
@@
Recently, Large language models (LLMs) have become powerful learners adept at handling multiple tasks, demonstrating remarkable reasoning capabilities and generalization across various domains, including natural language processing (NLP) [14], healthcare [15], law [16], and finance [17]. The GPT-4 has shown outstanding performance in NLP tasks, while TTM [18] has also demonstrated remarkable few-shot and zero-shot learning abilities in time series processing tasks. Inspired by these successes, there is growing interest in leveraging pre-trained models for cross-domain tasks, including channel-associated tasks in wireless communication. For instance, an LLM-empowered channel prediction method LLM4CP was proposed in [19], achieving greatly improved few-shot generalization ability. A foundational channel model WiFo [20] was trained on a diverse channel dataset to perform tasks like time-domain and frequency-domain prediction with zero-shot learning, but they primarily focus on channel reconstruction. Our objective is to enhance multiple channel-associated tasks in wireless communication using LLMs, which poses challenges such as effectively transferring cross-domain knowledge and managing task diversity. We address these challenges through the MoE-LoRA fine-tuning method [21], designed for multi-task scenarios, while noting that recent work has not sufficiently considered the specific relationships between channel-associated tasks, which may limit the model’s capacity to acquire general representations.
Specifically, we propose a wireless multi-task fine-tuning framework leveraging pre-trained large language models, which is successfully applied to the channel-associated multi-tasks. Unlike previous multi-task learning approaches that rely on shared bottom layers, we freeze most of the large model’s parameters and introduce MoE-LoRA for wireless multi-task fine-tuning. On the one hand, tasks can share experts’ weight, which helps the network learn common knowledge across tasks, while on the other hand, the independence of experts and the gating mechanism ensure the differentiation of task-specific features. Additionally, we incorporate multi-task adapter modules at both the input and output layers of the LLM to ensure alignment between the feature space of communication tasks and the semantic space of the pre-trained LLM. We also design a pre-processing method and corresponding output header for each task to enhance our framework further, optimizing the model’s overall performance and adaptability. The core contributions of this paper are summarized below:
@@
  • The model exhibits excellent performance on a range of wireless communication tasks, including channel estimation, channel prediction, localization enhancement, and beam management. Furthermore, it exhibits impressive generalization capabilities, highlighting robustness and versatility for diverse applications in the wireless domain.
@@ II SYSTEM DESCRIPTION
A dual-frequency communication system operating at sub-6G and mmWave frequencies is considered, which consists of one base station (BS) and one user equipment (UE). Both the BS and UE are equipped with transceivers for both frequency bands, with multiple-input multiple-output (MISO) and orthogonal frequency division multiplexing (OFDM) technologies applied to each. The sub-6G and mmWave antennas are co-located and aligned, with similar apertures, enabling them to share spatial features. For clarity, we use $\tilde{\left(\cdot\right)}$ notation, such as $\tilde{x}$ to denote parameters related to the sub-6G system. In mmWave band, the BS adopts an analog beamforming architecture and features $N_{t}$ antennas arranged in a uniform linear array (ULA) configuration, while in the sub-6G band, it utilizes a fully digital beamforming architecture with $\tilde{N}_{t}$ ULA antennas. The UE is equipped with an omnidirectional antenna for both frequency links and can accommodate multiple antennas through parallel processing.
@@ II-A Channel Model
In this context, $N$ and $P_{n}$ are the number of clusters and paths in each cluster, respectively. $\beta_{n,p}$, $\upsilon_{n,p}$, $\tau_{n,p}$, and $\Phi_{n,p}$ represent the complex path gain, doppler frequency shift, delay, and random phase, respectively. $\bm{a}(\theta_{n,p})$ represents the steering vector of the corresponding path, where $\theta_{n,p}$ denote the azimuth angles of departure (AoD). Considering the structural characteristics of ULA, the expression for $\bm{a}(\theta_{n,p})$ is derived as:
@@ II-B Signal Model
where $\bm{w}_{t}\in\mathbb{R}^{{N}_{v}\times 1}$ is the beam vector which is selected from predefined codebooks ${\bm{\mathcal{W}}}$, i.e. $\bm{w}_{t}\in{\bm{\mathcal{W}}}$. And $\bm{n}_{t,k}$ is the additive white gaussian noise (AWGN). The achievable spectral efficiency (SE) [22] of the downlink transmission process is derived as:
@@
Similarly, the sub-6G downlink signal transmission process also adopts MISO-OFDM, and it has the same transmission expression as the mmWave link, which will be omitted here for brevity. However, the sub-6G link employs digital precoding, so the uplink channel at the pilot positions is estimated using the Least Squares (LS) method as follows:
@@
where $\tilde{w}_{t,k}^{*}$ represents the beam vectors of time $t$ and the $-k\mbox{-}$th subcarrier at sub-6G link. Notably, the effectiveness of $\tilde{w}_{t,k}^{*}$ depends on the accuracy of $\tilde{{h}}_{LS}(t,f_{k})$. Specifically, inaccurate $\tilde{{h}}_{LS}(t,f_{k})$ will lead to mismatched $\tilde{w}_{t,k}^{*}$ with the actual $\tilde{h}(t,f_{k})$, reducing the received signal-to-noise ratio (SNR) and thereby impairing the SE. Therefore, accurate CSI is vital for improving the SE of the system.
@@ III-A Channel Reconstruction: Interpolation and Prediction
Channel reconstruction tasks typically aim to use known channel matrices to predict or interpolate the target channel matrices. The model is expected to capture the inter-domain correlations of the channel matrix across time, frequency, and antenna domains. So the channel matrix $\bm{H}$ considered herein includes three dimensions: time, space, and antenna, and can be expressed as follows:
@@
For the channel estimation task, a comb-type pilot pattern is used, featuring continuous pilots in the time-domain resource blocks (RBs) and a discrete arrangement in the frequency-domain RBs. The network is tasked with learning the channel characteristics across various frequency points and interpolating the missing channels at the absent frequency points. Consequently, $\bm{X}^{I}_{CE}$ and $\bm{X}^{L}_{CE}$ are defined as follows:
@@
Here, $\tilde{T}$ denotes the total number of timestamps, and $\tilde{K}$ denotes the number of OFDM symbols. $n_{pilot}$ denotes the number of pilots, which is typically set such that $\tilde{K}/n_{pilot}=4$.
@@ III-B Beam Management: Sub-6G Aided Beamforming
Beamforming requires accurate acquisition of the optimal weight vector $\mathbf{w}_{t}$ based on the codebook $\mathbf{W}\in\mathbb{R}^{N_{v}\times N_{c}}$, where $N_{c}$ is the codebook size. To achieve higher spatial resolution, a super-resolution Discrete Fourier Transform (DFT) codebook is applied, which implies that $N_{c}>N_{v}$. Since the main paths in the sub-6G and mmWave bands are often highly correlated under Line-of-Sight (LoS) conditions, leveraging sub-6 GHz to assist mmWave beamforming is highly beneficial, while also reducing the pilot overhead for mmWave beamforming. The $I_{BF}$ and $L_{BF}$ for sub-6G aided beamforming task can be defined as:
@@ III-C Radio Environment Mining: Distance Estimation and Path Loss Estimation
Due to the abundant environmental information present in the acquired channel data, such as the multipath components reflecting the density of surrounding buildings, and the delay indicating the distance to the target, the task of radio environment mining aims to leverage the estimated channel to extract environmental information, such as the distance $x_{d}$ from the UE to the BS and the path loss $x_{pl}$ of the main path. This extracted information can be utilized to adjust the configuration of the communication system, ultimately improving communication quality. Thus, $\bm{X}^{I}_{DE}$ and $\bm{X}^{L}_{DE}$ for distance estimation task can be defined as:
@@ IV LLM FOR Wireless Channel-Associated Tasks
All of these tasks mentioned above are fundamentally rooted in the wireless channel, suggesting that wireless multi-task learning can significantly enhance the model’s capacity to extract channel representations that generalize well. This, in turn, would improve the performance of each individual task. To leverage this potential, we propose a multi-task method for wireless channel applications, empowered by LLM, which we refer to as LLM4WM. This approach integrates the learning of these tasks within a unified network framework based on LLM. Below, a thorough description of the network components and the training process for LLM4WM are provided.
@@ IV-A Preprocessor Module
where the $\text{Flatten}(\bm{X},i)$ operation denotes the process of flattening the $i$-th dimension of the tensor $\bm{X}$, along with all subsequent dimensions, into a single dimension. For tasks such as beamforming, distance estimation, and path loss estimation that require channel angle features, the CSI data will undergo domain transformation to convert the spatial domain CSI to angle domain CSI, i.e.,
@@ IV-B Multi-Task Adapter Module
We extend the conventional use of adapter modules, which incorporate a small number of trainable parameters, enabling the model to preserve its existing modeling and generalization capabilities while adapting to a specific domain [28, 29]. However, it is primarily designed for single-task scenarios and lacks the ability to facilitate transfer generalization across multiple tasks simultaneously.
@@
where $L$ and $D_{\text{llm}}$ denote the token length of LLM’s input and LLM’s hidden dimension, respectively. Since the preprocessed features are two-dimensional data, the $\text{Linear}(\cdot)$ operation includes at least two fully connected operations, which linearly map the first and second dimensions to the specified dimensions. Then the residual feature extraction networks and the activation function would act on $\bm{X}_{n}^{f}$ to obtain feature maps with aligned semantic features as:
@@ IV-C Mixture-of-LoRA Based Fine-tuning
Assume the pre-trained weights are $W_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, where $d_{\text{in}}$ is the input dimension and $d_{\text{out}}$ is the output dimension. $A\in\mathbb{R}^{r\times d_{\text{in}}}$ and $B\in\mathbb{R}^{d_{\text{out}}\times r}$ are two trainable low-rank matrices. Then, we can obtain the fine-tune weights $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ as follows:
@@
where $r$ denotes the rank of the low-rank approximation. The hyperparameter $\alpha$ facilitates the adjustment of the rank $r$ and is typically set as $\alpha=2\times r$.
Assuming the input to the feed-forward network is $x_{t}$ and the output is $y_{t}$, the forward propagation of the model can be expressed as follows:
$y_{t}$ $=Wx_{t}=W_{0}x_{t}+\frac{\alpha}{r}BAx_{t}.$ (23)
@@
$y_{t}=Wx_{t}=W_{0}x_{t}+\frac{\alpha}{r}\sum_{k=1}^{N_{e}}{\omega}_{k}B_{k}A_{k}x_{t},$ (24)
@@
Notably, the design of the gating network has a direct impact on the performance of the MoE model. To prevent overfitting, we employ a single-layer linear network to generate the expert weights for each task and normalize the weight matrix using the $\text{Softmax}(\cdot)$ function to maintain the stability of the output data.
We apply MoE-LoRA to the linear layers within the feedforward network (FFN) of LLM while leaving the remaining parameters frozen. This approach significantly reduces the model’s trainable parameters, greatly lowering training costs and improving training efficiency.
@@ IV-D Multi-Task Output Module
To align the task’s output feature vector with the semantic space of the LLM, we use a multi-task adapter connected directly to the LLM’s output, which is identical to the input adapter. Assuming the task $n$’s output feature of the large model is $\bm{X}_{n}^{LLM}$, and the multi-task adapter for the task $n$ is denoted as $\text{Adapter}_{n}^{out}$, so this process can be represented as: